OpenStudy (anonymous):

Graph the line with slope -3 and passing through (2, 2).

OpenStudy (austinl):

Ok, we are given everything that we could possibly want for this problem! Slope intercept form is, \(y=mx+b\) Where m is slope. We can plug that in to the equation with what we know. \(y=-3x+b\) We are also given a point that it passes through, we would want to plug in those values into the equation. \((x,y)\rightarrow(2,2)\) \((2)=-3(2)+b\) We want to solve that for b, which is our y-intercept. \(b=2+6=8\) So that means that we have the equation, \(\color{blue}{y=-3x+8}\) Do you think you could graph that line? :D

OpenStudy (anonymous):

I think you are fantastic! Thanks so much. :)

OpenStudy (austinl):

You are more than welcome!